# Adaptive Rank, Reduced Forgetting: Knowledge Retention in Continual Learning Vision-Language Models with Dynamic Rank-Selective LoRA

### v3

@@ Abstract
We investigate whether the pre-trained knowledge of vision-language models (VLMs), such as CLIP, can be retained or even enhanced during continual learning (CL) while absorbing knowledge from a data stream. Existing methods often rely on additional reference data, isolated components for distribution or domain predictions, leading to high training costs, increased inference complexity, and limited improvement potential for pre-trained models. To address these challenges, we first comprehensively analyze the effects of parameter update locations and ranks on downstream adaptation and knowledge retention. Based on these insights, we propose Dynamic Rank-Selective Low Rank Adaptation (LoRA), a universal and efficient CL approach that adaptively assigns ranks to LoRA modules based on their relevance to the current data. Unlike prior methods, our approach continually enhances the pre-trained VLM by retaining both the pre-trained knowledge and the knowledge acquired during CL. Our approach eliminates the need for explicit domain or distribution prediction and additional reference data, enabling seamless integration of new tasks while preserving pre-trained capabilities. It also maintains the original architecture and deployment pipeline of the pre-trained model without incurring any additional inference overhead. Extensive experiments and analyses demonstrate that our method outperforms state-of-the-art approaches in continually absorbing knowledge of downstream tasks while retaining pre-trained knowledge.
@@ 1 Introduction
Figure 1: Illustrations of approaches for CL of VLMs. (a) Continual full fine-tuning is applied toward a universal CL updating. Additional reference data is used to alleviate the forgetting of pre-trained knowledge (zero-shot prediction on unseen data), e.g., [62]. It is computationally expensive and suffers from forgetting. (b) New task-specific components isolated to the pre-trained model are added in CL, e.g., [57, 54]. They require task/domain prediction/gating in inference, increasing inference complexity and restricting the pre-trained model’s improvement potential with limited usage scenarios. (c) We aim to continually learn the model by memory-efficient updating of all parameters in a universal manner, similar to (a), but without requiring reference data. It allows for enhancing the pre-trained model while learning continuously from the data stream, surpassing the limitations of (b). By utilizing dynamic rank-selective LoRA, we reduce forgetting of both the knowledge in the pre-trained model and the knowledge acquired during continual learning.
@@
Continual Learning (CL) [16, 9, 47, 25, 13] focuses on incrementally learning from new data streams without catastrophic forgetting [36, 32] or the need to retrain from scratch. With the advances of large-scale pre-trained models (PTMs), recent CL advances are increasingly leveraging these models. While many PTM-based CL approaches address continual adaptation for downstream tasks (e.g., class-incremental learning [51, 46, 20]), with an emphasis on downstream tasks’ performance, there is growing interest in preserving the inherent capabilities of PTMs, such as multi-domain continual learning for Vision-Language Models (VLMs) [62, 57, 54], or in continual pre-training to enhance PTMs over time [41]. In this paper, we focus on multi-domain continual learning for VLMs like CLIP [39].
@@
The ZSCL [62] framework aims to continually improve pre-trained model by incrementally feeding it with new knowledge. However, this often results in a degradation of both pre-trained and previously acquired knowledge during continual learning steps. To address this challenge, prior works have explored strategies like using reference data [62] or memory replay [25, 51, 40] to distill knowledge during full fine-tuning. However, these methods incur high training costs due to the need for additional data and extensive retraining (Fig. 1(a)), and yet they still suffered from the forgetting problem. Other approaches [57, 54] introduce task-specific components isolated from the pre-trained model during the CL stream, which require task/domain prediction or gating at inference. This increases inference complexity and limits the model’s adaptability to unseen data, restricting its generalization potential (Fig. 1(b)).
In this work, we explore a more general and practical form of CL for VLMs (similar to [62]), which maintains strong performance on seen data with continually updated pre-trained model, and retains or even boosting the performance on unseen data. To ensure memory-efficient learning, we utilize low-rank parameter updates (LoRA) [19]. Rather than inserting LoRA modules at fixed locations [26, 15], we extend them to all pre-trained weight matrices to enhance adaptation, and crucially, to analyze the impact of each weight matrix on downstream learning and knowledge retention.
Our analysis (Sec. 3.2) reveals that not all parameter updates contribute equally to downstream adaptation, with some potentially harming pre-trained capabilities. Manually selecting update locations and ranks can significantly improve the balance between adapting to new tasks and retaining pre-trained knowledge. This insight motivates our proposed method, Dynamic Rank-Selective LoRA, which adaptively determines both the locations within the pre-trained weight matrices and the ranks of parameter updates.
We leverage a reformulation of parameter updates using Singular Value Decomposition (SVD) [60, 12, 28, 35, 59]. Our method introduces learnable importance weights that mimic the singular values in SVD, capturing the significance of each rank in the parameter update. These learnable weights enable the model to dynamically adjust the contribution of each rank during training. To ensure only the most important updates are integrated into the pre-trained weights, we optimize the importance weights using a sparse prompting approach with a soft-thresholding operation. Our contributions are fourfold:
@@
    In this paper, we explore the universal CL of VLMs in an memory-efficient manner, maintaining strong performance on both seen and unseen data with reduced forgetting. We perform dynamic rank selective parameter updates adaptively to all pre-trained weight matrices.
@@
    Through comprehensive analyses, we show that the location and rank of parameter updates are critical for achieving an good balance between downstream task adaptation and pre-trained knowledge retention.
@@
    We propose a general low-rank parameter update method that adaptively perform dynamic rank selective updates, effectively integrating new knowledge while maximizing the retention of pre-trained knowledge.
@@
    Extensive experiments and in-depth analyses validate our approach across a wide range of benchmarks for downstream adaptation and zero-shot transfer. Additionally, we evaluate on hold-out data to further evaluate knowledge retention.
2 Related Works
@@ 3.1 Preliminaries
Continual Learning of VLMs. We focus on continual learning for VLMs within two challenging settings: multi-domain task-incremental learning (MTIL) [62, 57] and cross-domain task-agnostic incremental learning (X-TAIL) [54]. In the MTIL setting, the model sequentially learns from a set of $T$ tasks. For each task $t$ , the dataset is represented as $\mathcal{D}^{t}=\{({\mathbf{x}}^{t}_{i},y^{t}_{i})\}_{i=1}^{N^{t}}$, where ${\mathbf{x}}^{t}_{i}\in\mathbb{R}^{H\times W\times C}$ is the input image, $y^{t}_{i}\in\mathcal{C}^{t}$ is its corresponding class label, and $N^{t}$ is the number of samples in the task. The category set $\mathcal{C}^{t}=\{y^{t}_{j}\}_{j=1}^{M^{t}}$ consists of the class names within the task $t$, containing a total of $M^{t}$ classes.
In the X-TAIL setting, however, the category set during testing encompasses both previously seen and unseen domains, denoted as $\mathcal{C}=\mathcal{C}_{\text{seen}}\cup\mathcal{C}_{\text{unseen}}$. Here, $\mathcal{C}_{\text{seen}}=\bigcup_{i=1}^{n}\mathcal{C}_{i}$ includes all classes learned during prior tasks, while $\mathcal{C}_{\text{unseen}}$ represents novel classes that were not encountered during training. This introduces a more challenging scenario, as the model must classify test images across a diverse range of domains, making it more complex than the MTIL setting.
Transformers. A typical transformer model is composed of stacked blocks, where each block consists of two main submodules: an attention (Attn) module utilizing multi-head attention (MHA) and a multi-layer perceptron (MLP). In each transformer layer, the attention module at the $l$-th transformer layer can be defined as: $\text{Attn}(x)=\text{MHA}(x\mathbf{W}^{\text{Q}}_{l},x\mathbf{W}^{\text{K}}_{l},x\mathbf{W}^{\text{V}}_{l})\mathbf{W}^{\text{O}}_{l},$ where $\text{MHA}(\cdot)$ denotes the multi-head attention operation. The input sequence is denoted as $x\in\mathbb{R}^{n\times d}$, where $n$ is the sequence length and $d$ is the hidden dimension. The matrices $\mathbf{W}^{\text{Q}}_{l},\mathbf{W}^{\text{K}}_{l},\mathbf{W}^{\text{V}}_{l},\mathbf{W}^{\text{O}}_{l}\in\mathbb{R}^{d\times d}$ correspond to the Query, Key, Value, and Output projection matrices, respectively.
@@
Vision-Language Models (VLMs). In vision-language models like CLIP [39], the image encoder $f_{\theta}$ and text encoder $g_{\psi}$ are implemented as transformers in our case, respectively. Given an input image ${\mathbf{x}}\in\mathbb{R}^{H\times W\times C}$, it is segmented into patches and processed by the vision encoder to produce latent visual features ${\mathbf{z}}^{\text{V}}=f_{\theta}({\mathbf{x}})\in\mathbb{R}^{D}$. For text, the class label $y$ is embedded into prompts like “a photo of a [CLS]”, where [CLS] is the class name, resulting in text input ${\mathbf{t}}$. This prompt is encoded by the text encoder, yielding ${\mathbf{z}}^{\text{T}}=g_{\psi}({\mathbf{t}})\in\mathbb{R}^{D}$. During inference, the probability of classifying image ${\mathbf{x}}$ into class $y_{i}\in\{1,\dots,C\}$ is computed as: $p(y_{i}|{\mathbf{x}})=\frac{\exp(\mathtt{sim}({\mathbf{z}}^{\text{V}},{\mathbf{z}}^{\text{T}}{y_{i}})/\tau)}{\sum_{c=1}^{C}\exp(\mathtt{sim}({\mathbf{z}}^{\text{V}},{\mathbf{z}}^{\text{T}}_{y_{c}})/\tau)},$ where $\mathtt{sim}(\cdot)$ is cosine similarity, and $\tau$ is the temperature.
Figure 2: Evaluations of adaptation to downstream tasks and retention of zero-shot capabilities after training CLIP with various LoRA insertion locations and ranks.
Figure 3: Analysis of the effects of removing trained LoRA modules on downstream task adaptation (Aircraft) and retention of pre-trained capabilities (ImageNet-1k).
@@ 3.2 Analyses on LoRA Location and Rank
Analyses of locations and ranks of low-rank updates. We begin by investigating the influence of performing parameter updates at different locations within the pre-trained model, varying the update strength to assess its effects on both downstream adaptation and retention of pre-trained capabilities. The results of this analysis are presented in Fig. 3. We leverage Low-Rank Adaptation (LoRA) [19], where low-rank parameter updates are directly applied to the pre-trained weight matrices.
Our findings reveal the following: (1) Updating only the vision encoder leads to catastrophic forgetting of the pre-trained capabilities, significantly diminishing zero-shot performance. (2) Updating only the attention modules generally reduces forgetting but restricts the model’s ability to adapt to new tasks. (3) Using a higher rank enhances adaptation to downstream tasks without necessarily increasing the forgetting of pre-trained knowledge.
These observations highlight the need for a strategic selection of insertion locations and ranks to achieve an better balance between adaptation and zero-shot retention.
Analyses of pruning subsets of low-rank updates. To further investigate, we analyzed the effects of selectively perturbing individual modules of the pre-trained weights, as shown in Fig. 3. Initially, we inserted LoRA into all pre-trained matrices with a rank of 16, corresponding to the “Vision + Text, Attn + MLP, rank 16” configuration shown in Fig. 3. We then iteratively remove the LoRA updates learned for each module, examining the impact on both downstream adaptation and the knowledge retention.
Our findings reveal that not all parameter updates are equally critical for downstream adaptation, and some may even degrade the model’s pre-trained capabilities. For example, removing updates in the Value projection of the Attention module within the vision encoder had minimal impact on downstream adaptation but significantly restored the model’s pre-trained zero-shot performance. Similarly, removing updates in the MLP layer of the vision encoder slightly reduced downstream task performance but, interestingly, enhanced the model’s pre-trained capabilities, even surpassing its original zero-shot performance.
3.3 CL VLMs with Dynamic Rank-Selective LoRA
@@
Overview. As shown in Fig. 4, we introduce dynamic rank-selective parameter updates for all pre-trained weight matrices, adaptively inserting updates based on the importance of each rank for current training data of task $t$. Specifically, we integrate dynamic rank-selective LoRA into all pre-trained weights within the Attention and MLP modules of both the vision and text encoders. For clarity, for learning on task $t$, we denote the pre-trained weights or those updated from the previous task as $\{\mathbf{W}^{t,m}_{0}\}_{m=1}^{M}$, where $M$ represents the total number of pre-trained weight matrices added with LoRA. After training on each task $t$, these dynamic rank updates are merged back into the original pre-trained weights.
@@
Rank-selective LoRA with importance weights. LoRA [19] is a parameter-efficient fine-tuning technique designed to adapt pre-trained models using a minimal number of additional parameters. Instead of directly updating a pre-trained weight matrix $\mathbf{W}_{0}^{t,m}\in\mathbb{R}^{d\times k}$, LoRA introduces two low-rank matrices, $\mathbf{B}^{t,m}\in\mathbb{R}^{d\times r}$ and $\mathbf{A}^{t,m}\in\mathbb{R}^{r\times k}$, such that $\Delta\mathbf{W}^{t,m}=\mathbf{B}^{t,m}\mathbf{A}^{t,m}$, where $r\ll\min(d,k)$. The updated weight matrix is defined as:
In this formulation, the original weights $\mathbf{W}_{0}^{t,m}$ remain fixed, and only the matrices $\mathbf{B}^{t,m}$ and $\mathbf{A}^{t,m}$ are trained, significantly reducing the number of trainable parameters to $r(d+k)$. Notably, during inference, LoRA incurs no additional computational overhead compared to the original pre-trained model, as the low-rank updates can be merged into the pre-trained weights, as shown in Eq. (1).
Previous works reformulate the low-rank parameter update in the form of Singular Value Decomposition (SVD) [60, 12, 28, 35, 59]. We introduce dynamic importance weights into the formulation:
$\Delta\mathbf{W}^{t,m}=\sum_{i=1}^{r}\mathbf{w}_{i}^{t,m}\mathbf{B}_{i}^{t,m}\mathbf{A}_{i}^{t,m},$ (2)
where $\mathbf{w}^{t,m}\in\mathbb{R}^{r}$ is a learnable vector that represents the importance of each rank during training. The importance weights imitate the diagonal elements in the singular value matrix of SVD and we allow these weights to be learned directly from the data.
The introduction of these importance weights incurs only a minimal increase in trainable parameters, adding just $r$ additional parameters per LoRA module. For initialization, we adhere to the convention that randomly initialize the matrix $\mathbf{A}$, set the matrix $\mathbf{B}$ to zero, and we initialize the importance weights $\mathbf{w}^{t,m}$ randomly.
Dynamic rank selection with sparse importance. The formulation in Eq. (2) enables the model to dynamically adjust the importance of each rank through gradient descent. [^1] To align the training process with the pre-trained model, we use CLIP’s contrastive loss during pre-training as the supervised training loss. However, not all ranks contribute equally to downstream task performance (as shown in Fig. 3). To ensure that only the most impactful LoRA ranks are retained in the pre-trained weights, we propose to optimize the learning of importance weights in a sparse-promoting way by minimizing the $\ell_{1}$ norm of the importance weights. The optimization objective for paramters $\{\mathbf{w}^{t,m},\mathbf{B}^{t,m},\mathbf{A}^{t,m}\}_{m=1}^{M}$ learned on task $t$ are as follows:
$\mathcal{L}_{\text{train}}^{t}:=\mathcal{L}_{\text{sup}}^{t}+\lambda\sum_{m=1}^{M}\|\mathbf{w}^{t,m}\|_{1},$ (3)
where $\mathcal{L}_{\text{sup}}^{t}$ is the supervised training loss, $M$ denotes the total number of pre-trained weight matrices inserted with LoRA of rank $r$, and $\mathbf{w}^{t,m}$ represents the importance weights of the each rank of the LoRA module added to the pre-trained weight matrix $\mathbf{W}^{t,m}_{0}$, and $\lambda$ controls the strength of the $\ell_{1}$ regularization.
We adopt the proximal gradient [4] for handling the non-differentiable $\ell_{1}$ regularization applied to the importance weights. Specifically, the dynamic update of the importance weight for the $i$-th rank of LoRA can be written as a soft-thresholding operation:
@@
where $\hat{\mathbf{w}}^{t,m}_{i}$ denotes the value of ${\mathbf{w}^{t,m}_{i}}$ after gradient update using only the supervised loss $\mathcal{L}_{\text{sup}}^{t}$. The threshold parameter $\kappa$ is initially set to zero and gradually increases until it reaches a maximum value $\kappa_{\text{max}}$. The indicator function $\mathbbm{1}(\cdot)$ returns 1 if condition is met, 0 otherwise, and $\mathrm{sign}(\cdot)$ represent the +/+/-+ / - sign of input.
This encourages that only the ranks with significant importance are preserved, while those with low importance are pruned. By adjusting the weightings iteratively, the model focuses dynamic parameter updates on the most relevant components. In the early stages of training, dense updates are performed without Eq. (4) to allow all ranks to capture task-relevant information.
Continually enhance the pre-trained model. After training is complete, ranks with importance weight of zero are removed, while non-zero ranks are merged into the pre-trained weights by substituting Eq. (2) into the original LoRA formulation Eq. (1). When a new task is introduced, a new set of LoRA modules is initialized, and the computations involve only the newly initialized LoRA modules in conjunction with the continually updated pre-trained model. This approach allows the model to leverage accumulated knowledge efficiently while keeping the computational footprint minimal without additional overhead.

### v4

@@ Abstract
We investigate whether the pre-trained knowledge in vision-language models (VLMs), such as CLIP, can be retained—or even enhanced—in continual learning (CL) while incorporating new knowledge from the data stream. Existing CL methods primarily focus on continual downstream adaptation using components isolated from pre-trained model (PTM), increasing inference complexity and limiting improvements to the PTM itself; some also retain knowledge relying on additional reference data, leading to high training costs. To address these limitations, we propose a universal and efficient Continual Learning approach for VLM based on Dynamic Rank-Selective LoRA (CoDyRA), which directly improves the PTMs while preserving the existing knowledge from both pre-training and CL. Through analyses on how LoRA rank and placement impact/regularize the learning and forgetting in CL, we propose CoDyRA that adaptively performs rank-minimized parameter updates in different modules, based on their importance to the current data. This ensures a balance between knowledge acquisition (plasticity) and forgetting mitigation (stability). Our method operates without explicit domain or distribution prediction and does not rely on reference data, enabling seamless task integration while maintaining pre-trained capabilities. Moreover, CoDyRA preserves the original model architecture and deployment pipeline, introducing no additional inference overhead. Extensive experiments demonstrate that our approach enhances representations based on new downstream data while retaining pre-trained knowledge, achieving state-of-the-art results.
@@ 1 Introduction
Figure 1: Illustrations of different approaches for CL in VLMs. (a) Continual full fine-tuning is applied to achieve universal CL updates, often incorporating additional reference data to mitigate the forgetting of pre-trained knowledge (e.g., zero-shot prediction on unseen data, as in [63]). It is computationally expensive and suffers from forgetting. (b) New task-specific components isolated from the PTM are added for CL, e.g., [58, 55]. They require task/domain prediction/gating in inference, increasing inference complexity and restricting the PTM’s improvement potential with restricted usage scenarios. (c) We aim to continually train the model via parameter-and-data-efficient adaptation in a universal manner, without requiring reference data or full fine-tuning as in (a). It enhances the PTM while continually learning from the data stream, surpassing the limitations of (b) on unseen data.
@@
Continual learning (CL) [17, 10, 48, 27, 14] focuses on incrementally learning from new data streams without catastrophic forgetting [37, 34] or the need to retrain with all seen data. With the advances of large-scale pre-trained models (PTMs), recent studies increasingly perform CL with these models [66, 35, 47, 22, 55]. Many PTM-based CL approaches primarily focus on continual adaptation for downstream tasks (e.g., class-incremental learning, CIL [43, 52, 22, 35]), with an emphasis on downstream task performance. While task-specific designs [66, 35, 55] offer promising improvements in adaptation performance, preserving the inherent capabilities of PTMs during continual learning is crucial for maintaining and enhancing generalizable representations. This has garnered growing interest in research areas such as multi-domain continual learning for VLMs [63], or in continual pre-training to enhance PTMs over time [42]. In this paper, we focus on CL of VLMs like CLIP [40] with multi-domain data.
@@
Some approaches, such as ZSCL [63], aim to continually update PTMs by incrementally incorporating new data with full fine-tuning while mitigating forgetting of pre-trained knowledge through memory replay [27, 52, 41] using additional reference data. However, these methods incur high training costs due to the reliance on additional data and extensive retraining (Fig. 1(a)), yet they still struggle with forgetting. To retain PTM capabilities during CL on downstream tasks, some approaches [58, 55] introduce task-specific components that isolated from the PTM during the CL stream, which require task/domain prediction or gating at inference. This increases inference complexity and limits the model’s adaptability to unseen data, restricting its generalization potential (Fig. 1(b)).
In this work, we propose a practical and efficient CL approach for VLMs that continually trains and updates pre-trained models without relying on downstream task-specific components. Our method preserves and even enhances the generalizable representations of PTMs while seamlessly integrating new knowledge from data streams. To achieve efficient and effective learning, we conduct comprehensive analyses of low-rank adaptation (LoRA) [20] for CL in PTMs. Rather than restricting LoRA to fixed locations as in prior works [28, 16], we extend it to all parameter weights for a thorough investigation, aiming to approximate the capabilities of full fine-tuning. We analyze how the rank and placement of LoRA updates impact or regularize the learning dynamics, influencing both downstream adaptation and the retention of existing knowledge (Sec. 3.2).
Our analyses reveal that not all LoRA ranks and placements contribute equally to downstream learning and forgetting. Generally, for arbitrary weights, relatively higher-rank LoRA facilitates learning new tasks but tends to increase forgetting, while lower-rank LoRA mitigates forgetting but limits adaptation. [^1] Similar findings are also observed in the language domain [5]. However, we identify that there usually exists a balance between plasticity (learning new knowledge) and stability (regularizing forgetting) at a moderately small (but not too small) rank, potentially maximizing CL benefits. Moreover, we observe that this balance varies significantly across different parameter placements.
The analysis results motivate us to propose Continual Dynamic Rank-Selective LoRA (CoDyRA), which performs parameter updates via dynamically optimizing for a minimized LoRA rank suitable for the current data. The rank selection for each LoRA is achieved through a Singular Value Decomposition (SVD)-based reformulation [61, 13, 30, 36, 60], with learnable importance weights corresponding to the singular values. With sparsity-promoting regularization, these learnable weights allow the model to dynamically adjust the contribution of each rank during training. This ensures that only the most essential updates (i.e., LoRA with a minimized rank) are integrated into the pre-trained weights, enabling learning from new data while mitigating forgetting. Our contributions are fourfold:
@@
    We investigate a universal and practical CL paradigm for efficiently updating pre-trained VLMs on new data streams while mitigating forgetting, without relying on reference data or additional components.
@@
    We conduct in-depth analyses on the impact of LoRA ranks and placements, revealing their crucial role in balancing downstream learning and knowledge retention.
@@
    We propose CoDyRA, which dynamically optimizes a minimized rank for parameter updates based on current data, with learnable importance weights for each rank.
@@
    Extensive experiments across multiple CL benchmarks validate our method’s superior performance in both downstream learning and mitigating forgetting, and additionally prove superiority on pre-trained knowledge retention using hold-out data.
2 Related Work
@@ 3.1 Preliminaries
Continual Learning of VLMs. We focus on CL of VLMs with multiple domain data, e.g., multi-domain task-incremental learning (MTIL) [63, 58] and cross-domain task-agnostic incremental learning (X-TAIL) [55]. In the MTIL setting, the model sequentially learns from a set of $T$ tasks. For each task $t$ , the dataset is represented as $\mathcal{D}^{t}=\{({\mathbf{x}}^{t}_{i},y^{t}_{i})\}_{i=1}^{N^{t}}$, where ${\mathbf{x}}^{t}_{i}\in\mathbb{R}^{H\times W\times C}$ is the input image, $y^{t}_{i}\in\mathcal{C}^{t}$ is its corresponding class label, and $N^{t}$ is the number of samples in the task. The category set $\mathcal{C}^{t}=\{y^{t}_{j}\}_{j=1}^{M^{t}}$ represents the class labels for task $t$, consisting of $M^{t}$ classes in total. In the X-TAIL setting, the category set during testing includes classes from both previously seen and unseen domains, denoted as $\mathcal{C}=\mathcal{C}_{\text{seen}}\cup\mathcal{C}_{\text{unseen}}$. Here, $\mathcal{C}_{\text{seen}}=\bigcup_{i=1}^{n}\mathcal{C}_{i}$ includes all classes learned during prior tasks, while $\mathcal{C}_{\text{unseen}}$ represents novel classes that were not encountered during training. This poses a more challenging scenario, as the model must classify test images across a broader range of domains.
Transformers. A typical transformer model is composed of stacked blocks, where each block consists of two main submodules: an attention (Attn) module utilizing multi-head attention (MHA) and a multilayer perceptron (MLP). In each transformer layer, the attention module at the $l$-th layer is defined as: $\text{Attn}(x)=\text{MHA}(x\mathbf{W}^{\text{Q}}_{l},x\mathbf{W}^{\text{K}}_{l},x\mathbf{W}^{\text{V}}_{l})\mathbf{W}^{\text{O}}_{l},$ where $\text{MHA}(\cdot)$ denotes the multi-head attention operation. The input sequence is denoted as $x\in\mathbb{R}^{n\times d}$, where $n$ is the sequence length and $d$ is the hidden dimension. The matrices $\mathbf{W}^{\text{Q}}_{l},\mathbf{W}^{\text{K}}_{l},\mathbf{W}^{\text{V}}_{l},\mathbf{W}^{\text{O}}_{l}\in\mathbb{R}^{d\times d}$ correspond to the Query, Key, Value, and Output projection matrices, respectively.
@@
Vision-Language Models (VLMs). In vision-language models like CLIP [40], both the image encoder $f_{\theta}$ and text encoder $g_{\psi}$ are implemented as transformers in our case. Given an input image ${\mathbf{x}}\in\mathbb{R}^{H\times W\times C}$, it is segmented into patches and processed by the vision encoder to produce latent visual features ${\mathbf{z}}^{\text{V}}=f_{\theta}({\mathbf{x}})\in\mathbb{R}^{D}$. For text, the class label $y$ is embedded into prompts like “a photo of a [CLS]”, where [CLS] is the class name, resulting in text input ${\mathbf{t}}$. This prompt is encoded by the text encoder, yielding ${\mathbf{z}}^{\text{T}}=g_{\psi}({\mathbf{t}})\in\mathbb{R}^{D}$. During inference, the probability of classifying image ${\mathbf{x}}$ into class $y_{i}\in\{1,\dots,C\}$ is computed using the softmax function as: $p(y_{i}|{\mathbf{x}})=\frac{\exp(\mathtt{sim}({\mathbf{z}}^{\text{V}},{\mathbf{z}}^{\text{T}}{y_{i}})/\tau)}{\sum_{c=1}^{C}\exp(\mathtt{sim}({\mathbf{z}}^{\text{V}},{\mathbf{z}}^{\text{T}}_{y_{c}})/\tau)},$ where $\mathtt{sim}(\cdot)$ is cosine similarity, and $\tau$ is the temperature.
@@ 3.2 Analyses on LoRA Location and Rank
To achieve efficient learning, we explore LoRA and aim to optimize its usage strategy. To analyze the impact of LoRA’s rank and placement on downstream learning and forgetting, we conduct a series of experiments: (1) We apply LoRA components with varying ranks at different locations in the pre-trained CLIP model (Fig. 4); (2) We apply LoRA across all PTM weights and systematically prune them for evaluation (Fig. 4). We train the model on downstream datasets and assess its performance on both (1) the new task and (2) an additional reference dataset to evaluate the preservation of pre-trained capabilities.
Applying LoRA on all weights. As shown in Fig. 4 and Fig. 4, our analysis of LoRA placement shows that applying LoRA to specific modules alone may limit learning capacity in different ways: (1) The different LoRA placement locations significantly affect downstream performance and the forgetting/retention of pre-trained capabilities; (2) Updating only the vision encoder leads to more forgetting of pre-trained capabilities; (3) Updating only the attention modules helps mitigate forgetting but limits the learning ability on new tasks. These observations align with the distinct roles of different PTM components. Rather than manually selecting LoRA placement, we apply it across all weights and optimize for an adaptive configuration.
The effects of different LoRA ranks on balancing learning and forgetting. From Fig. 4, we can observe that, despite learning low-rank space, a relatively higher-rank LoRA facilitates learning new tasks but tends to increase forgetting, while lower-rank LoRA mitigates forgetting but limits adaptation. This suggests that a lower-rank LoRA produces stronger regularization in maintaining the learned knowledge by sacrificing the learning of new knowledge. Importantly, we identify a balance at a moderately small rank, which reduces forgetting while maintaining adaptability. This implies that minimizing the rank while fitting on new data may find an optimized trade-off for CL.
The balance varies by location and task. The ranks for a good balance vary by placement (Fig. 4), and pruning LoRA at different locations has varying effects (Fig. 4), indicating that not all LoRA components contributed equally for a task. For example, removing updates in the Value projection of the Attention module in the vision encoder minimally affected adaptation but significantly restored pre-trained performance. Interestingly, pruning the MLP layer of the vision encoder slightly reduced task performance but enhanced pre-trained capabilities. These findings motivate our adaptive approach to dynamically determine ranks based on location and task.
Figure 3: Evaluations of downstream task adaptation and zero-shot capability retention after training CLIP with different LoRA insertion locations and ranks. Different shapes represent encoders, colors indicate transformer modules, and sizes denote ranks.
Figure 4: Analysis of the impact of removing trained LoRA modules on downstream task adaptation (Aircraft) and retention of pre-trained capabilities (ImageNet-1k).
3.3 Proposed CoDyRA
Overview. Our analyses in Sec. 3.2 suggest applying LoRA across all CLIP modules while adaptively selecting an appropriate rank for each task and location to balance learning and forgetting. While finding an ideally global optimal rank is challenging, we approach it by jointly optimizing downstream task performance and minimizing the number of ranks using sparsity-promoting regularization.
@@ 3.3 Proposed CoDyRA
As shown in Fig. 2, we introduce dynamic rank-selective parameter updates for all pre-trained weight matrices within the Attention and MLP modules of both the vision and text encoders. These updates are adaptively inserted based on the importance of each rank for the current training data. For clarity, during training on task $t$, we denote the pre-trained weights—or those updated from the previous tasks—as $\{\mathbf{W}^{t,m}_{0}\}_{m=1}^{M}$, where $M$ represents the total number of pre-trained weight matrices augmented with LoRA. After training on each task $t$, these dynamic rank updates are merged back into the original pre-trained weights.
Low-rank updates with rank importance. For efficient adaptation, instead of directly updating a pre-trained weight matrix $\mathbf{W}_{0}^{t,m}\in\mathbb{R}^{d\times k}$, LoRA [20] introduces two low-rank matrices, $\mathbf{B}^{t,m}\in\mathbb{R}^{d\times r}$ and $\mathbf{A}^{t,m}\in\mathbb{R}^{r\times k}$, such that $\Delta\mathbf{W}=\mathbf{B}^{t,m}\mathbf{A}^{t,m}$, where $r\ll\min(d,k)$. The updated weight matrix is then defined as:
In this formulation, the original weights $\mathbf{W}_{0}^{t,m}$ remain fixed, and only $\mathbf{B}^{t,m}$ and $\mathbf{A}^{t,m}$ are trained, reducing the number of trainable parameters to $r(d+k)$. Notably, LoRA incurs no additional computational overhead during inference, as the low-rank updates can be efficiently merged into the pre-trained weights, as shown in Eq. (1).
To automatically and adaptively decide the rank in the LoRA updates, we introduce a learnable importance weight vector $\mathbf{w}^{t,m}\in\mathbb{R}^{r}$ and incorporate them into the formulation of LoRA as follows:
$\Delta\mathbf{W}^{t,m}=\sum\nolimits_{i=1}^{r}\mathbf{w}_{i}^{t,m}\mathbf{B}_{i}^{t,m}\mathbf{A}_{i}^{t,m},$ (2)
where $\mathbf{w}^{t,m}$ is learnable and represents the importance of each rank. The formulation is based on a reformulation of LoRA updates in Eq. (1) based on Singular Value Decomposition (SVD) [61, 13, 30, 36, 60]. The importance weights are corresponding to the diagonal elements in the singular value matrix of SVD.
Introducing importance weights incurs minimal overhead, adding only $r$ additional parameters per module. For initialization, we follow the standard convention: matrix $\mathbf{A}$ is randomly initialized, matrix $\mathbf{B}$ is set to zero, and the importance weights $\mathbf{w}$ are randomly initialized.
@@
Dynamic rank-selective updates with sparse rank importance. The formulation in Eq. (2) enables the model to dynamically adjust the importance of each rank through gradient descent. [^2] To align the training process with the pre-trained model, we use CLIP’s contrastive loss as the optimization objective during pre-training. However, not all ranks contribute equally to downstream task performance (as shown in Fig. 4). To ensure that only the most impactful ranks are retained for learning on new task while removing unnecessary ranks to mitigate forgetting of already-learned knowledge, we propose optimizing the importance weights with a sparsity-promoting regularization through minimizing the $\ell_{1}$ norm. The optimization objective for parameters $\{\mathbf{w}^{t,m},\mathbf{B}^{t,m},\mathbf{A}^{t,m}\}_{m=1}^{M}$ learned on task $t$ is:
$\mathcal{L}_{\text{train}}^{t}:=\mathcal{L}_{\text{sup}}^{t}+\lambda\sum\nolimits_{m=1}^{M}\|\mathbf{w}^{t,m}\|_{1},$ (3)
where $\mathcal{L}_{\text{sup}}^{t}$ is the supervised training loss, $M$ is the total number of pre-trained weight matrices inserted $r$-rank update, and $\mathbf{w}^{t,m}$ represents the importance weights of each rank added to the weight matrix $\mathbf{W}^{t,m}_{0}$. The $\ell_{1}$ regularization strength is controlled by $\lambda$.
To handle the non-differentiable $\ell_{1}$ regularization applied to the importance weights, we adopt the proximal gradient method [4]. The $i$-th element of $\hat{\mathbf{w}}^{t,m}$ for rank $i$ is updated via soft-thresholding:
@@
where $\hat{\mathbf{w}}^{t,m}_{i}$ denotes the value of ${\mathbf{w}^{t,m}_{i}}$ after applying the gradient update from the supervised loss $\mathcal{L}_{\text{sup}}^{t}$. The threshold parameter $\kappa$ is initially set to zero and gradually increases to a maximum value $\kappa_{\text{max}}$, analogous to linearly scheduling the regularization strength $\lambda$. The indicator function $\mathbbm{1}(\cdot)$ returns 1 if the condition is met and 0 otherwise, while $\mathrm{sign}(\cdot)$ represents the +/+/-+ / - sign of the input.
This approach preserves only the ranks with significant importance, pruning those with low relevance. By iteratively adjusting weightings, the model focuses dynamic parameter updates on the most relevant components. During the early stages of training, dense updates are performed without Eq. (4) to allow all ranks to capture task-relevant information before applying sparsity constraints.
Continually enhancing the pre-trained model. After training, ranks with zero importance weights are pruned, while non-zero ranks are merged into the pre-trained weights by substituting Eq. (2) into the original LoRA formulation Eq. (1). When a new task is introduced, a fresh set of CoDyRA modules is initialized, and computations involve only these newly initialized CoDyRA modules alongside the continually updated pre-trained model. This approach enables the model to efficiently leverage accumulated knowledge while maintaining a minimal computational footprint, ensuring no additional overhead.